OpenStudy (anonymous):

Evaluate the integral integral of sec^3(2x)dx I have tried it so many times and I keep getting it wrong. My answer is: 1/4(tan(2x)sec(2x)-log(cos(x)-sin(x))+log(sin(x)+cos(x)))

hartnn (hartnn):

your answer is correct. just add +c to it.

OpenStudy (anonymous):

oh i forgot to add that in the open question. but, it still says wrong.

OpenStudy (jamesj):

Well, start with integration by parts, \[ \int \sec^2(2x) \sec(2x) \ dx = \frac{1}{2} \tan(2x) \sec(2x) - \int \tan^2(2x)\sec(2x) \ dx \] Now, that integral on the right hand side equals \[ \int \sec^3(2x) \ dx - \int \sec(2x) \ dx \] Thus rearranging \[ 2 \int \sec^3(2x) \ dx = \frac{1}{2} \tan(2x)\sec(2x) + \int \sec(2x) \ dx \] So far so good?

hartnn (hartnn):

then maybe simplify the 2 log terms using logA-log B = log (A/B)

OpenStudy (anonymous):

yupp good.

hartnn (hartnn):

log [ (cos + sin)/(cos -sin) ]

hartnn (hartnn):

we can avoid integration by parts(if wanted to), if we can use some standard results....

OpenStudy (jamesj):

Now, you know how to integrate sec(2x)? Multiply by (sec 2x + tan 2x) / (sec 2x + tan 2x) and we have \[ \int \frac{ \sec^2(2x) + \sec(2x)\tan(2x)}{\tan 2x + \sec 2x} dx = \frac{1}{2}\ln( \tan 2x + \sec 2x) \] Now you pull everything together

OpenStudy (anonymous):

I got it! It was a simple problem. I forgot a multiplication sign lol!

OpenStudy (jamesj):

ok